OpenStudy (twizttiez):

The system of equations is coincident. What are the missing values? -2x+5y=-12 -6x+___y=___?

OpenStudy (twizttiez):

I think it's -6x+(-2)y=3

OpenStudy (twizttiez):

Can you help @Catlover5925 ?

OpenStudy (catlover5925):

i have no idea, why do you think it is -2 and 3

OpenStudy (twizttiez):

@Miracrown

OpenStudy (twizttiez):

One moment

OpenStudy (catlover5925):

@wio

OpenStudy (twizttiez):

-2x+5y=-12 eqn 1 -6x+___y=___? eqn 2 the eqn 2 is 3 times the eqn 1 3(-2x + 5y) = 3(-12)

OpenStudy (catlover5925):

that sounds right @geerky42 and @wio please check this

OpenStudy (twizttiez):

I do need help on one more would either of you two mind helping?

geerky42 (geerky42):

coincident means both equation is exactly the same. http://www.emathematics.net/posicionrelativa.php So you have \(-2x+5y=-12\) Multiply both sides by 3 and you have \(-6x+15y=-36\) Hence missing values would be 15 and -36.

OpenStudy (twizttiez):

Thanks so much could you help on one more?

OpenStudy (twizttiez):

Column B A. consistent independent B. coincident C. inconsistent